OpenStudy (anonymous):

The regression equation y = 0.5x + 1.3 approximates the number of hours it takes to arrange a dinner event, y, given the number of people attending, x. Which statement is true? For every extra hour taken to make arrangements for an event, the attendance increases by 0.5 people. For every extra hour taken to make arrangements for an event, the attendance increases by 1.3 people. For every extra person attending an event, the time it takes to make arrangements increases by 0.5 hour. For every extra person attending an event, the time it takes to make arrangements increases by 1.3 hours.

OpenStudy (alyssajobug):

For every extra person attending an event, the time it takes to make arrangements increases by 0.5 hour.

OpenStudy (alyssajobug):

say 50 people attend and plug it in y=0.5(50)+1.3 y=25+1.2 y=26.2

OpenStudy (alyssajobug):

Now change it to 100 people y=0.5(100)+1.3 y=50+1.3 y=51.3

OpenStudy (alyssajobug):

Notice, it goes up as you add the amount of people by 0.5

OpenStudy (anonymous):

The regression equation y = –0.414x + 106.55 approximates the percent of people in an audience who finish watching a documentary, y, given the length of the film in minutes, x. Which is the best prediction for the percent of people in an audience who will finish watching a documentary that is 180 minutes long? 75% 67% 32% 26%

OpenStudy (alyssajobug):

y=-0.414(180)+106.55

OpenStudy (alyssajobug):

y=74.52+106.55

OpenStudy (alyssajobug):

y=181.07

OpenStudy (alyssajobug):

You see, in each of these cases you needed to see that the x value is given and plug it in to find the value of y

OpenStudy (alyssajobug):

I have to go to bed now but feel free to ask me any questions and I will answer them when I am on in the morning, simply tag me by adding @alyssajobug to your question and have a nice night!

OpenStudy (anonymous):

Thanks :)